OpenStudy (anonymous):

Find the fifth roots of 243(cos 300° + i sin 300°).

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

@SithsAndGiggles Can you help me?

OpenStudy (anonymous):

Do you mean \[\sqrt[5]{243\left(\cos300^\circ+i\sin300^\circ\right)}~~?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well, \(243=3^5\), and using Euler's formula, you know that \(\cos300^\circ+i\sin300^\circ=e^{300^\circ i}\). So, the above number is equivalently \[3\sqrt[5]{e^{300 i}}=3\left(e^{300 i}\right)^{1/5}=3e^{60 i}\] Applying Euler's formula again, you have \[3\left(\cos60^\circ+i\sin60^\circ\right)\] Simplify that and you get your answer.

OpenStudy (anonymous):

I don't think this is the approach that I'm supposed to be taking :/ ... For example, for this example problem that I found it says: Find the fifth roots of 32(cos 280° + i sin 280°). and the answers are: 2*[ cos(56) + isin(56) ] 2*[ cos(128) + isin(128) ] 2*[ cos(200) + isin(200) ] 2*[ cos(272) + isin(272) ] 2*[ cos(344) + isin(344) ]

OpenStudy (anonymous):

Using the same method as for the last one: \[\sqrt[5]{32\left(\cos280^\circ+i\sin280^\circ\right)}=2\left(e^{280i}\right)^{1/5}=2e^{56i}=2\left(\cos56^\circ+i\sin56^\circ\right)\]

OpenStudy (anonymous):

And that list implies a multiple choice question, since none of the other options are equal to each other.

OpenStudy (anonymous):

But it's not a multiple choice problem :/

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

In that case, I'm clearly missing some information that only you have access to. I don't know how you may have been taught this material; I only did what I thought was right for this problem. Sorry I couldn't help.

OpenStudy (anonymous):

It could also be the case that you didn't simplify your answer. When you have \(3\left(\cos60^\circ+i\sin60^\circ\right)\), you can simplify the sine and cosine: \[\cos60^\circ=\frac{1}{2}\\ \sin60^\circ=\frac{\sqrt3}{2}\]